OpenStudy (anonymous):

Given the polynomial 6x3 + 4x2 - 6x - 4, what is the value of the coefficient 'k' in the factored form? 6x3 + 4x2 - 6x - 4 = 2(x + k)(x - k)(3x + 2)

OpenStudy (butterflydreamer):

Firstly we know that 6x^3 + 4x^2 - 6x - 4 = 2 (3x + 2 ) (x + k) (x - k) Note: i just re-arranged your question but it's exactly the same so don't freak out ^_^ Anyways, So basically from your question it means that when we multiply 2 * (3x + 2) * (x + k) * (x - k) we will get 6x^3 + 4x^2 - 6x - 4 So using what we have, we want to focus on this part.  What i mean is we need to find out what this part equals to " 2 ( 3x+ 2)" = ? Can you tell me what 2 (3x + 2) equals to? :)